Assume that the risk-free rate is 3.5% and the market risk premium is 5%.

- What is the required return for the overall stock market? Round your answer to two decimal places.

- What is the required rate of return on a stock with a beta of 0.6? Round your answer to two decimal places.
Answer #1

1. They are 2 methods to calculate it.
a. the required return for the overall stock market = market risk premium= 5%.

b. Capital Asset pricing model:

As per CAPM model:
Re= Rf+(Rm-Rf)B

Re= required rate of return of the stock market. ?
Rf= Risk-free rate. 3.5%
Rm =Market Risk Premium. 5%
B = Beta, systematic risk. The overall stock market has a Beta of 1.

Re = 3.5+(5-3.5)1
Re= 5%

the required return for the overall stock market = 5%

2. Capital Asset pricing model:

As per CAPM model:
Re= Rf+(Rm-Rf)B

Re= required rate of return. ?
Rf= Risk-free rate (nominal) = 3.5%
Rm =Market Risk Premium. 5%
B = Beta, systematic risk. 0.6

Re= 3.5+(5-3.5) x 0.6
Re= 4.4%

the required rate of return on a stock with a beta of 0.6 = 4.4%